OpenStudy (anonymous):

Factoring Trinomials with Negative Numbers x2 + x – 12 6. x2 – 2x – 8

OpenStudy (anonymous):

If you FOIL (x+4)(x-3) what do you get?

OpenStudy (anonymous):

ok thnk

OpenStudy (anonymous):

u got it? great!

OpenStudy (anonymous):

#6 is kinda hard in a way

OpenStudy (anonymous):

you need factors of -8 that sum to -2

OpenStudy (anonymous):

-1*8 1*-8 -2*4 2*-4

OpenStudy (anonymous):

which pair adds up to -2?

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

yeah im so lost

OpenStudy (anonymous):

Here is the factored version of #6: (x+2)(x-4)

OpenStudy (anonymous):

F O I L watch what happens when I FOIL it: x^2 -4x +2x -8

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

now simplify the middle two terms =x^2 -2x -8 notice how the factors of 8=-4*2 add up to the middle term -2? then look at the factored form and you see the -4 and +2 in the factors

OpenStudy (anonymous):

good?